. Sea water contains an average of 1.08 x 103 ppm of Na+ and 270 ppm of SO4-2. Calculate 

a. The molar concentration of Na+and SO4-2 given that the average density  of sea water is 1.02 g/mL 

b. The pNa and pSO4 for sea water 




1
Expert's answer
2021-09-07T02:08:02-0400

Given: ppm of Na+=1.08×103+=1.08×103 ppm. The density of sea water = 1.02 g/mL.

molar concentration of Na+ Na+ = 1.08 * 10^3 ppm = 1080 mg/L = 1080 * 10^-3g/23g/mol = 0.046M SO42- = 270 * 10^-3/96 = 2.81 * 10^-3

= 4.79×10"^{-2}" mol/L4.79×10-2mol/L